Shutdown karlsuss2 SNF 2009-05-07 09:06:31: FSE working on BSA

gsosa at snf.stanford.edu gsosa at snf.stanford.edu
Thu May 7 17:19:32 PDT 2009


Unable to focus on mask with BSA microscopes. Checked and reseated all connectors associated with BSA microscopes. Checked voltages on distribution board. Voltage OK at + and - 10VDC referenced to ground. Checked voltage on power supply in DVM unit. All OK- +/- 5 VDC and +/- 12 VDC. Rechecked BSA focus. Now able to focus on mask with BSA microscopes. FSE verified setup with BSA setup mask- All OK.. No adjustment needed. FSE checked WEC level- OK, no adjustment needed. FSE checked WEC spring force tension. Tension too high at 1200 newtons. Adjusted WEC pressure for 850 newtons. PLEASE do not adjust WEC pressure. Vendor is still working on alignment issues and will retur tomorrow. BSA OK to use TSA OK to use . Please continue to follow instructions posted on tool for fine alignmnet.




More information about the karlsuss2-pcs mailing list